I am floored. For the love of sweet Jesus someone please help me.

The question is to try and find the largest possible area of a right triangle that has a 28 cm long hypotenuse.

Units are square centimeters.

I've tried a really hard to figure it out but I keep getting it wrong. Please help!
Answer #1

The area of a right triangle of legs a,b is A = \frac{1}{2}ab

The hypotenuse is c2 = a2 + b2

We also know that if a,b are both real value then the square of is difference is non-negative: (a−b)2≥0. (Strictly positive unless a=b.)

If we solve: (a−b)2 = a2+2ab+b2,  where 2ab=4A. So we can replace and

0\leq{(a-b)^{2}}

0 \leq {a^{2}}+2ab+b^{2}

0 \leq {c^{2}}-4A

4A\leq {c^{2}}

A\leq\frac{1}{4}c^{2}

So we already have this identity: A\leq\frac{1}{4}c^{2}

In this problem c=28 so our answer is:

So our answer is:

A\leq\frac{1}{4}28^{2} = \frac{1}{2} * 784 = 196

A\leq196

Largest possible area = 196 square centimetres.